If Mavis is 5 yrs older than her brother. 5 years ago she was twice her brother's age, how old are they now?
Found 2 solutions by oberobic, stanbon:
Answer by oberobic(2304)

M = Mavis' age
B = Brother's age
.
M = B+5
.
Five years ago...
M-5
B-5
.
Mavis was twice her brother's age.
(M-5) = 2(B-5)
.
Substitute M = B+5
.
(B+5-5) = 2B-10
.
B = 2B-10
-B = -10
B = 10
.
M = B+5 = 15
.
Check their ages 5 years ago.
B = 10-5 = 5
M = 15-5 = 10
M = 2*B
Correct.
.
Answer: Mavis is 15 and her brother is 10 years old.
.
Done.

Answer by stanbon(75887)

If Mavis is 5 yrs older than her brother. 5 years ago she was twice her brother's age, how old are they now?
---------------------
Equations:
m = b + 5
m-5 = 2(b-5)
-------
Substitute for "m" and solve for "b":
b+5-5 = 2b - 10
brother = 10 yrs (Brother's age now)
--------
Solve for "m":
m = b + 5
m = 10+5 = 15 yrs (Mavis's age now)
====================================
Cheers,
Stan H.
==============